# Global Convergence of Policy Gradient for Entropy Regularized Linear-Quadratic Control with Multiplicative Noises

### v3

@@ Appendix C Proofs in Section 4
$=\mu\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\left[\frac{1}{1-\gamma}-\frac{1}{1-\gamma V_{K}}\right]$
@@
$S_{K}$ $=\mathbb{E}[x_{0}^{2}]\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{\infty}\gamma^{t}(1-V_{K}^{t})$ (18)
@@
$S^{(l)}_{K,\Sigma}$ $=\mu\sum_{t=0}^{l-1}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{l-1}\gamma^{t}(1-V_{K}^{t})$ (19)
@@
$\leq(Q+K^{\top}RK)\gamma^{l}S_{K,\Sigma}+(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)\right]\gamma^{l}$ (20)
$=\gamma^{l}\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]$ (21)
@@ C.3 Two useful Lemmas
$|(1-\gamma V_{K})^{-1}-(1-\gamma V_{K^{\prime}})^{-1}|$ $\leq 2(1-\gamma V_{K})^{-2}|\gamma V_{K^{\prime}}-\gamma V_{K}|$
@@
$\leq 2(1-\gamma V_{K})^{-1}|\gamma V_{K^{\prime}}-\gamma V_{K}|$
@@
We have $\|\Sigma^{\prime}\|\leq 2\|\Sigma\|$ as $\|\Sigma^{\prime}\|-\|\Sigma\|\leq\|\Sigma^{\prime}-\Sigma\|\leq\|\Sigma\|$
@@
$:=\Delta_{\Sigma}(1-\gamma V_{K})^{-1}$
@@
$:=h_{K}\|K^{\prime}-K\|+h_{2}\|\Sigma-\Sigma^{\prime}\|$
@@
$\leq(E_{K^{\prime}}-E_{K})S_{K^{\prime},\Sigma^{\prime}}+(S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma})E_{K}$
@@
$\|\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)\|$ $\leq\|E_{K^{\prime}}-E_{K}\||S_{K^{\prime},\Sigma^{\prime}}|+|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|E_{K}$
@@
$=\left[\|R\|+\gamma A\cdot h_{5}\|B\|+\gamma P_{K}(B^{\top}B+D^{\top}D)++2\gamma\cdot h_{5}\|B^{\top}B+D^{\top}D\|\|K\|\right]\|K^{\prime}-K\|$
@@
$\|E_{K^{\prime}}-E_{K}\|S_{K^{\prime},\Sigma^{\prime}}\leq(S_{K,\Sigma}+h_{K}h_{\Sigma}+h_{2}\|\Sigma\|)h_{E}\|K^{\prime}-K\|$
@@
$|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|\|E_{K}\|\leq(h_{K}\|K^{\prime}-K\|+h_{2}\|\Sigma^{\prime}-\Sigma\|)\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}$ (25)
@@
$+\left(h_{2}\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}\right)\|\Sigma^{\prime}-\Sigma\|$
@@
Apply theorem 35 in [5] we have
@@
if $\|\Sigma-\Sigma^{\prime}\|\leq\frac{\sigma_{min}(\Sigma)}{2}$
@@
$=\frac{\gamma\left(P_{K^{\prime}}-P_{K}\right)(B^{\top}B+D^{\top}D)-\frac{\tau}{2}(\Sigma^{\prime-1}-\Sigma^{-1})}{(1-\gamma)}$
@@
$:=h_{8}\|K^{\prime}-K\|+h_{9}\|\Sigma^{\prime}-\Sigma\|_{F}$
@@
$\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)=(\nabla_{K}f_{r_{1}}(K,\Sigma)-\nabla_{K}f(K,\Sigma))+(\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma))$
@@
$\leq\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$
@@
$\quad\leq\sigma_{1}:=\big(\frac{2n}{r_{1}}f(K,\Sigma)\big)^{2}+\big(\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}\big)^{2}$
@@
$\widehat{\nabla}_{K}-\nabla_{K}f(K,\Sigma)=(\widehat{\nabla}_{K}-\nabla_{K}^{\prime})+(\nabla_{K}^{\prime}-\widetilde{\nabla}_{K})+(\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma))$
@@
$\|\nabla_{K}^{\prime}\|\leq\frac{\epsilon}{3}+\|\widetilde{\nabla}_{K}\|\leq\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$
@@
$:=2\Gamma L^{2}f(K,\Sigma)r_{1}+\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$
Define $Z_{i}:=\big[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\big]U_{i}$
@@
$:=(2\Gamma L^{2}f(K,\Sigma)r_{1})^{2}+(\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|})^{2}$
@@
$\mathbb{P}\left(\|\widehat{\nabla}_{K}-\nabla_{K}^{\prime}\|\leq\frac{\epsilon}{3}\right)\leq 1-\sqrt{\kappa_{1}}$
@@
$\mathbb{P}\left(\|\widehat{\nabla}_{K}-\nabla_{K}f(K,\Sigma)\|\leq{\epsilon}\right)\leq 1-\kappa_{1}$
@@
$\widehat{\nabla}_{L}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}|x_{t})\right)\right]V_{i}$
@@
$\widetilde{\nabla}_{L}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}$
@@
$\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)$ $=(\nabla_{L}f_{r_{2}}(K,L)-\nabla_{L}f(K,\Sigma))+(\widetilde{\nabla}_{L}-\nabla_{L}f_{r_{2}}(K,L))$ (27)
@@
$\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}\|\leq\|\frac{2n(n+1)}{2r_{2}^{2}}f(K,\Sigma)V_{i}\|\leq\frac{n(n+1)}{r_{2}}f(K,\Sigma)$
@@
$\|\nabla_{L}f_{r_{2}}(K,L)\|_{F}\leq\|\nabla_{L}f(K,\Sigma)\|_{F}+\frac{\epsilon}{6}$
$\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}-\nabla_{L}f_{r_{2}}(K,L)\|_{F}\leq\frac{n(n+1)}{r_{2}}f(K,\Sigma)+\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{\epsilon}{6}$
@@
$\quad\leq\sigma_{Z_{1}}:=\frac{n(n+1)}{r_{2}}f(K,\Sigma)$
@@
$\widehat{\nabla}_{L}-\nabla_{L}f(K,\Sigma)=(\widehat{\nabla}_{L}-\nabla_{L}^{\prime})+(\nabla_{L}^{\prime}-\widetilde{\nabla}_{L})+(\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma))$
For the third term, based on the previous proof, we have $\|\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with probability $1-\kappa_{2}$ under conditions given in the lemma. Furthermore, we have $\|\widetilde{\nabla}_{L}\|\leq\|\nabla_{L}f(K,\Sigma)\|+\frac{\epsilon}{3}$
@@
$\leq\frac{n(n+1)}{2r_{2}}\frac{1}{M}\sum_{i=1}^{M}\left(f^{(l)}(K,\Sigma_{i})-f(K,\Sigma_{i})\right)$
If set
$l\geq\frac{\log\frac{2r_{2}\epsilon}{3n(n+1)}-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{log\gamma}$
@@
$\|\nabla_{L}^{\prime}\|\leq\|\widetilde{\nabla}_{L}\|+\frac{\epsilon}{3}\leq\|\nabla_{L}f(K,\Sigma)\|+\frac{2\epsilon}{3}\leq\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{2\epsilon}{3}$
@@
$\leq R_{Z}:=\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)+\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{2\epsilon}{3}$
@@
$\quad\leq\sigma_{Z_{2}}:=\big(\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)\big)^{2}$
@@
$=2L\nabla_{\Sigma}f(K,\Sigma)$
@@
with probability $1-\kappa$
@@
$\widehat{S}-{S_{K,\Sigma}}=\widehat{S}-\widetilde{S}^{(l)}+\widetilde{S}^{(l)}-\widetilde{S}+\widetilde{S}-\ {S_{K,\Sigma}}$
@@
$|q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma}|\leq|q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma^{\prime}}|+|q_{K,\Sigma^{\prime}}-q_{K,\Sigma}|$
@@
$:=h_{11}\|\Sigma^{\prime}-\Sigma\|_{F}$

### v4

@@ Appendix C Proofs in Section 4
$=\mu\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\left[\frac{1}{1-\gamma}-\frac{1}{1-\gamma V_{K}}\right].$
@@
$S_{K}$ $=\mathbb{E}[x_{0}^{2}]\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{\infty}\gamma^{t}(1-V_{K}^{t}),$ (16)
@@
$S^{(l)}_{K,\Sigma}$ $=\mu\sum_{t=0}^{l-1}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{l-1}\gamma^{t}(1-V_{K}^{t}).$ (17)
@@
$\leq(Q+K^{\top}RK)\gamma^{l}S_{K,\Sigma}+(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)\right]\gamma^{l}$
$=\gamma^{l}\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right].$
@@ C.3 Two useful Lemmas
$|(1-\gamma V_{K})^{-1}-(1-\gamma V_{K^{\prime}})^{-1}|$ $\leq 2(1-\gamma V_{K})^{-2}|\gamma V_{K^{\prime}}-\gamma V_{K}|.$
@@
$\leq 2(1-\gamma V_{K})^{-1}|\gamma V_{K^{\prime}}-\gamma V_{K}|.$
@@
We have $\|\Sigma^{\prime}\|\leq 2\|\Sigma\|$ as $\|\Sigma^{\prime}\|-\|\Sigma\|\leq\|\Sigma^{\prime}-\Sigma\|\leq\|\Sigma\|.$
@@
$:=\Delta_{\Sigma}(1-\gamma V_{K})^{-1}.$
@@
$:=h_{K}\|K^{\prime}-K\|+h_{2}\|\Sigma-\Sigma^{\prime}\|.$
@@
$\leq(E_{K^{\prime}}-E_{K})S_{K^{\prime},\Sigma^{\prime}}+(S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma})E_{K}.$
@@
$\|\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)\|$ $\leq\|E_{K^{\prime}}-E_{K}\||S_{K^{\prime},\Sigma^{\prime}}|+|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|E_{K}.$
@@
$=\left[\|R\|+\gamma A\cdot h_{5}\|B\|+\gamma P_{K}(B^{\top}B+D^{\top}D)++2\gamma\cdot h_{5}\|B^{\top}B+D^{\top}D\|\|K\|\right]\|K^{\prime}-K\|.$
@@
$\|E_{K^{\prime}}-E_{K}\|S_{K^{\prime},\Sigma^{\prime}}\leq(S_{K,\Sigma}+h_{K}h_{\Sigma}+h_{2}\|\Sigma\|)h_{E}\|K^{\prime}-K\|,$
@@
$|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|\|E_{K}\|\leq(h_{K}\|K^{\prime}-K\|+h_{2}\|\Sigma^{\prime}-\Sigma\|)\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|},$
@@
$+\left(h_{2}\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}\right)\|\Sigma^{\prime}-\Sigma\|.$
@@
Apply Theorem 3.5 in [5] we have
@@
if $\|\Sigma-\Sigma^{\prime}\|\leq\frac{\sigma_{min}(\Sigma)}{2}$.
@@
$=\frac{\gamma\left(P_{K^{\prime}}-P_{K}\right)(B^{\top}B+D^{\top}D)-\frac{\tau}{2}(\Sigma^{\prime-1}-\Sigma^{-1})}{(1-\gamma)}.$
@@
$:=h_{8}\|K^{\prime}-K\|+h_{9}\|\Sigma^{\prime}-\Sigma\|_{F}.$
@@
$\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)=(\nabla_{K}f_{r_{1}}(K,\Sigma)-\nabla_{K}f(K,\Sigma))+(\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)).$
@@
$\leq\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}.$
@@
$\quad\leq\sigma_{1}:=\big(\frac{2n}{r_{1}}f(K,\Sigma)\big)^{2}+\big(\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}\big)^{2}.$
@@
$\widehat{\nabla}_{K}-\nabla_{K}f(K,\Sigma)=(\widehat{\nabla}_{K}-\nabla_{K}^{\prime})+(\nabla_{K}^{\prime}-\widetilde{\nabla}_{K})+(\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)).$
@@
$\|\nabla_{K}^{\prime}\|\leq\frac{\epsilon}{3}+\|\widetilde{\nabla}_{K}\|\leq\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|},$
@@
$:=2\Gamma L^{2}f(K,\Sigma)r_{1}+\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}.$
Define $Z_{i}:=\big[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\big]U_{i}$.
@@
$:=(2\Gamma L^{2}f(K,\Sigma)r_{1})^{2}+(\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|})^{2}.$
@@
$\mathbb{P}\left(\|\widehat{\nabla}_{K}-\nabla_{K}^{\prime}\|\leq\frac{\epsilon}{3}\right)\leq 1-\sqrt{\kappa_{1}}.$
@@
$\mathbb{P}\left(\|\widehat{\nabla}_{K}-\nabla_{K}f(K,\Sigma)\|\leq{\epsilon}\right)\leq 1-\kappa_{1}.$
@@
$\widehat{\nabla}_{L}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}|x_{t})\right)\right]V_{i}.$
@@
$\widetilde{\nabla}_{L}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i},$
@@
$\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)$ $=(\nabla_{L}f_{r_{2}}(K,L)-\nabla_{L}f(K,\Sigma))+(\widetilde{\nabla}_{L}-\nabla_{L}f_{r_{2}}(K,L)).$ (22)
@@
$\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}\|\leq\|\frac{2n(n+1)}{2r_{2}^{2}}f(K,\Sigma)V_{i}\|\leq\frac{n(n+1)}{r_{2}}f(K,\Sigma),$
@@
$\|\nabla_{L}f_{r_{2}}(K,L)\|_{F}\leq\|\nabla_{L}f(K,\Sigma)\|_{F}+\frac{\epsilon}{6}.$
$\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}-\nabla_{L}f_{r_{2}}(K,L)\|_{F}\leq\frac{n(n+1)}{r_{2}}f(K,\Sigma)+\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{\epsilon}{6}.$
@@
$\quad\leq\sigma_{Z_{1}}:=\frac{n(n+1)}{r_{2}}f(K,\Sigma),$
@@
$\widehat{\nabla}_{L}-\nabla_{L}f(K,\Sigma)=(\widehat{\nabla}_{L}-\nabla_{L}^{\prime})+(\nabla_{L}^{\prime}-\widetilde{\nabla}_{L})+(\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)).$
For the third term, based on the previous proof, we have $\|\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with probability $1-\kappa_{2}$ under conditions given in the lemma. Furthermore, we have $\|\widetilde{\nabla}_{L}\|\leq\|\nabla_{L}f(K,\Sigma)\|+\frac{\epsilon}{3}.$
@@
$\leq\frac{n(n+1)}{2r_{2}}\frac{1}{M}\sum_{i=1}^{M}\left(f^{(l)}(K,\Sigma_{i})-f(K,\Sigma_{i})\right).$
If
$l\geq\frac{\log\frac{2r_{2}\epsilon}{3n(n+1)}-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{log\gamma},$
@@
$\|\nabla_{L}^{\prime}\|\leq\|\widetilde{\nabla}_{L}\|+\frac{\epsilon}{3}\leq\|\nabla_{L}f(K,\Sigma)\|+\frac{2\epsilon}{3}\leq\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{2\epsilon}{3}.$
@@
$\leq R_{Z}:=\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)+\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{2\epsilon}{3}.$
@@
$\quad\leq\sigma_{Z_{2}}:=\big(\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)\big)^{2}.$
@@
$=2L\nabla_{\Sigma}f(K,\Sigma),$
@@
with probability $1-\kappa$.
@@
$\widehat{S}-{S_{K,\Sigma}}=\widehat{S}-\widetilde{S}^{(l)}+\widetilde{S}^{(l)}-\widetilde{S}+\widetilde{S}-\ {S_{K,\Sigma}}.$
@@
$|q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma}|\leq|q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma^{\prime}}|+|q_{K,\Sigma^{\prime}}-q_{K,\Sigma}|.$
@@
$:=h_{11}\|\Sigma^{\prime}-\Sigma\|_{F}.$